OpenStudy (anonymous):

The rectangular pool at the Peabody is 27 feet wide. To put border title around the pool the contractor would need 134 feet of tile? How many feet long is the pool?

OpenStudy (anonymous):

So, you know that Perimeter = P = 2L + 2W, then P = 134, so: 134 = 2L + 2*27 134 - 54 = 2L + 54 -54 80 = 2L L = 40. So, the pool is 40 feet long.